# Mapping the Spiral of Silence: Surveying Unspoken Opinions in Online Communities

### v1

@@ A.1 Generating Controversial Viewpoints
Role of social media platforms in spreading fake news Social media’s lax approach has led to a fake news epidemic, undermining informed decision-making and threatening societal stability. Blaming social media for fake news is misplaced; users should fact-check and stricter content control risks censorship.
@@
Role of the United States in the Venezuelan political crisis I believe US intervention is vital for Venezuela’s democratic restoration and humanitarian aid. I view US involvement in Venezuela as neo-imperialism; each country should independently handle its internal affairs.
@@
Table 5: Our viewpoint generation process can be used for non-political communities. We present a sample of topics, and their respective viewpoints, proposed for r/NBA and r/gaming.
Topic Viewpoint 1 Viewpoint 2
@@
LeBron James vs Michael Jordan LeBron’s consistent dominance and achievements arguably make him the GOAT, surpassing MJ. Despite LeBron’s feats, MJ’s NBA Finals record, killer instinct, and game-changing ability make him the GOAT in my opinion.
NBA Season Length I believe the 82-game season should be shortened to maintain game quality and reduce player fatigue. The 82-game NBA season is crucial for testing team endurance, reducing fluke performances, and providing ample basketball for fans.
International NBA games I believe international NBA games disrupt the season’s rhythm, fatigue players with travel, and unfairly risk their health and team performance for sport’s geographical expansion. I support international NBA games as they expand the audience, grow the brand, inspire youth, and boost the sport’s global popularity, despite logistical challenges.
NBA Draft System I strongly favor abolishing the draft system for a free market system, allowing players to choose their teams and potentially balance league competition. The NBA draft is crucial for maintaining balance between small and large market teams and preventing the formation of super-teams.
Golden State Warriors 2015-2019 The Warriors’ 2015-2019 run boosted the league’s competitiveness by forcing teams to adapt and improve The Warriors’ 2015-2019 dominance made the NBA predictable and potentially drove fans away.
@@
Console vs PC As a console gamer, I value its simplicity, exclusives, and plug-and-play convenience for relaxation. As a PC enthusiast, I believe PCs provide unmatched customization, performance, and game variety for in-depth gaming.
@@
Working conditions in gaming Crunch culture in game development is harmful and unsustainable; companies should prioritize employee wellbeing. Crunch may not be ideal, but without it, could top games be made? Many vocations are stressful, not just gaming.
@@
Large vs indie game companies Major game companies’ dominance stifles creativity; we need more space for indie developers to innovate. Big companies like Nintendo prove that dominance doesn’t necessarily kill creativity, as they continually innovate and fund ambitious projects.
@@
eSports as a real sport eSports, despite lacking physical exertion, is a sport due to its demand for strategy, teamwork, and skill. Esports should be classified as competitive gaming, not traditional sports due to the lack of physical exertion.
@@
Objectification of female characters I believe game developers overly sexualize female characters, neglecting character depth. Video games are fantasy and idealized characters reflect artistic vision, not objectification.
@@
Copyright issues and modding I support modders as they express themselves by reshaping games, while respecting authorship. I support developers because unauthorized mods can threaten their control over their costly game creations.
A.2 Identifying Relevant Subreddits
@@
To find our list of political and social issues-oriented subreddits, we use the following process. Starting with 2,040 subreddits from r/ListOfSubreddits, we look at the 50 hottest submissions in each subreddit as of January 5, 2024. We select subreddits where the majority of submissions are in English – classified using langdetect (Shuyo 2010) – leaving us with 1,975 subreddits. Next, using GPT-3.5-Turbo we classify whether post titles are political (see below for prompt). We choose subreddits where more than $80\%$ of submissions are labelled as being political (N=48). From the remaining subreddits, we select our final list of 23 based on subreddit size. In our final survey, participants also had the option to list an additional subreddit they were an active member of, bringing us to 33 subreddits in total.
@@
1Political content on Reddit is varied and can be about officials and activists, social issues, or news and current events. Looking at the following post title, would you categorize it as POLITICAL or NOT POLITICAL content?
2
3Answer 1 if it is POLITICAL, 0 otherwise.
4
5Post: \${POST}
6Answer:
A.3 Knowledge Quiz
@@
To validate that participants were members of the subreddits they selected, they were given a knowledge screening quiz. In the quiz, the participants were shown three post titles, two of which belonged to top 50 hottest submissions of their selected subreddit and one that came from the top 50 hottest submissions of a different political subreddit. We manually checked that the title from the other subreddit was not relevant to the selected subreddit. The titles used in the knowledge quiz is listed in Table 6.
@@
Table 6: Examples of post titles used in the knowledge quizzes. Participants must correctly select the post titles that belong to the subreddit (i.e., those in columns “Title 1” and “Title 2”).
A.4 Survey Details
@@
A.5 Descriptive Statistics
Finally, we provide the descriptive statistics of the variables we use in our linear mixed-effect models in Table 7.
Table 7: An overview of descriptive statistics for all introduced variables (min, max, mean, median).
@@ A.5 Descriptive Statistics
Dependent Variables Share Likelihood 1 7 3.79 4
Upvote Likelihood 1 7 5.41 6
Controls WTSC 1.5 5.5 3.68 3.75
Agreement Intensity 1 3 2.13 2
Male (0/1) 0 1 0.70 1
White (0/1) 0 1 0.68 1
Democrat (0/1) 0 1 0.34 0
Account Tenure 0 16 5.80 5
Posting 1 4 2.51 2
@@
Inclusion 1 11 6.36 7
Diversity 1 11 6.29 6
Content Removal Rate (%) 4.13 46.04 29.77 34.47
@@ Appendix B Robustness Checks
In this section, we conduct three robustness checks on our analyses. First, we report results using an ordinal mixed effect regression model in Table 8. Second, we use an alternative measure (ModRatio) for moderation. In our pre-registration, we planned on using both ModRatio and Content Removal Rate as measures of community moderation but ultimately removed ModRatio as the two were correlated. Finally, we provide results for our analysis including a quadratic term for Content Removal Rate.
Ordinal mixed effect regression
In Table 8, we use an ordinal mixed effect regression to predict Share Likelihood, which is a Likert-scale with values ranging from $1$ to $7$. We find that this set of models yields qualitatively similar results to Models 1a-c reported in Sec 5. There is a statistically significant negative relationship between Incongruency ($\beta=-0.66$, $p<0.001$) and the likelihood of sharing a viewpoint. Furthermore, there is a positive relationship between the dependent variable and Inclusion ($\beta=0.28$, $p<0.01$) as well as a interaction effect between Inclusion and Incongruency ($\beta=0.37$, $p<0.05$).
Table 8: An ordinal mixed effect model yields qualitatively similar results to the linear mixed effect model reported in Sec. 5.
Model 3a Model 3b Model 3c
Coef. SE $p$ Coef. SE $p$ Coef. SE $p$
Dependent Variables (Share Likelihood)
Share Likelihood = 1 $-1.83$ $0.25$ $0.00^{***}$ $-2.24$ $0.29$ $0.00^{***}$ $-2.29$ $0.28$ $0.00^{***}$
Share Likelihood = 2 $-0.96$ $0.24$ $0.00^{***}$ $-1.33$ $0.28$ $0.00^{***}$ $-1.36$ $0.27$ $0.00^{***}$
Share Likelihood = 3 $-0.55$ $0.24$ $0.02^{*}$ $-0.91$ $0.28$ $0.00^{**}$ $-0.94$ $0.27$ $0.00^{***}$
Share Likelihood = 4 $-0.11$ $0.24$ $0.63$ $-0.44$ $0.27$ $0.11$ $-0.47$ $0.27$ $0.08$
Share Likelihood = 5 $0.84$ $0.24$ $0.00^{***}$ $0.58$ $0.27$ $0.03^{*}$ $0.55$ $0.26$ $0.04^{*}$
Share Likelihood = 6 $1.98$ $0.26$ $0.00^{***}$ $1.77$ $0.28$ $0.00^{***}$ $1.73$ $0.28$ $0.00^{***}$
@@
WTSC $-0.41$ $0.08$ $0.00^{***}$ $-0.40$ $0.08$ $0.00^{***}$ $-0.41$ $0.09$ $0.00^{***}$
Agreement Intensity $0.44$ $0.08$ $0.00^{***}$ $0.36$ $0.08$ $0.00^{***}$ $0.39$ $0.08$ $0.00^{***}$
Male (0/1) $-0.22$ $0.18$ $0.22$ $-0.31$ $0.18$ $0.09$ $-0.33$ $0.18$ $0.07$
White (0/1) $0.01$ $0.17$ $0.94$ $0.03$ $0.17$ $0.85$ $0.05$ $0.17$ $0.77$
Democrat (0/1) $-0.40$ $0.17$ $0.03^{*}$ $-0.64$ $0.18$ $0.00^{***}$ $-0.64$ $0.17$ $0.00^{***}$
Account Tenure $0.22$ $0.08$ $0.00^{***}$ $0.17$ $0.08$ $0.03^{*}$ $0.17$ $0.08$ $0.04^{*}$
Posting $0.46$ $0.08$ $0.00^{***}$ $0.37$ $0.08$ $0.00^{***}$ $0.40$ $0.08$ $0.00^{***}$
@@
(H1) Incongruency (0/1) $-0.73$ $0.18$ $0.00^{***}$ $-0.66$ $0.19$ $0.00^{***}$
(H2a) Inclusion $0.41$ $0.09$ $0.00^{***}$ $0.28$ $0.11$ $0.01^{**}$
(H2b) Diversity $0.04$ $0.09$ $0.56$ $0.04$ $0.10$ $0.68$
(H3) Content Removal Rate $0.25$ $0.15$ $0.10$ $0.27$ $0.15$ $0.07$
(H2a) Incongruency $\times$ Inclusion $0.37$ $0.19$ $0.05^{*}$
(H2b) Incongruency $\times$ Diversity $0.04$ $0.18$ $0.84$
(H3) Incongruency $\times$ Content Removal Rate $-0.22$ $0.19$ $0.24$
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Replacing measures for moderation
We examine changes to the model when replacing Content Removal Rate with ModRatio as our measure of moderation. ModRatio is the ratio of the number of subscribers to number of moderators in a subreddit. In this case, a higher ModRatio would mean less active moderation as there are fewer moderators per subscribers. We apply a logarithmic transformation with base 2 after adding a start-value of 1 to ModRatio. As shown in Table 9, we find a significant negative relationship between ModRatio and Share Likelihood ($p=0.04$). However, the impact on the dependent variable is minimal. When we look at model coefficients using non-standardized variables, a $1\%$ increase in the ratio of subscribers to moderators only leads to a $0.003$ increase in the likelihood of sharing an opinion. The trends with our other independent variables are the same as when we used Content Removal Rate.
Examining quadratic terms
We also report results of our linear mixed-effects model after including a quadratic term for Content Removal Rate in Table 10. Again, we see similar trends between Inclusion and Incongruency on Share Likelihood. There is no significant relationship between the quadratic term and our dependent variable.
Table 9: Including ModRatio does not substantially alter relationships between our independent variables and the likelihood of sharing a viewpoint. The results of a linear mixed-effect model that predicts a participant’s likelihood of sharing their opinion to a subreddit. We present results including ModRatio as an independent variable and results after adding interaction effects.

### v2

@@ A.2.2. Generated Topics and Viewpoints
Role of social media platforms in spreading fake news 1. Social media’s lax approach has led to a fake news epidemic, undermining informed decision-making and threatening societal stability.
2. Blaming social media for fake news is misplaced; users should fact-check and stricter content control risks censorship.
@@
Role of the United States in the Venezuelan political crisis 1. I believe US intervention is vital for Venezuela’s democratic restoration and humanitarian aid.
2. I view US involvement in Venezuela as neo-imperialism; each country should independently handle its internal affairs.
Overview of the eleven political topics and their opposing viewpoints used in the opinion expression survey. Overview of the eleven political topics and their opposing viewpoints used in the opinion expression survey. The table lists eleven political topics used. Each is paired with two contrasting statements designed to capture different sides of the debate.
@@
Table 6. Our viewpoint generation process can be used for non-political communities. We present a sample of topics, and their respective viewpoints, proposed for r/NBA and r/gaming.
Topic Viewpoint
@@
LeBron James vs Michael Jordan 1. LeBron’s consistent dominance and achievements arguably make him the GOAT, surpassing MJ.
2. Despite LeBron’s feats, MJ’s NBA Finals record, killer instinct, and game-changing ability make him the GOAT in my opinion.
NBA Season Length 1. I believe the 82-game season should be shortened to maintain game quality and reduce player fatigue.
2. The 82-game NBA season is crucial for testing team endurance, reducing fluke performances, and providing ample basketball for fans.
International NBA games 1. I believe international NBA games disrupt the season’s rhythm, fatigue players with travel, and unfairly risk their health and team performance for sport’s geographical expansion.
2. I support international NBA games as they expand the audience, grow the brand, inspire youth, and boost the sport’s global popularity, despite logistical challenges.
NBA Draft System 1. I strongly favor abolishing the draft system for a free market system, allowing players to choose their teams and potentially balance league competition.
2. The NBA draft is crucial for maintaining balance between small and large market teams and preventing the formation of super-teams.
Golden State Warriors 2015–2019 1. The Warriors’ 2015–2019 run boosted the league’s competitiveness by forcing teams to adapt and improve.
2. The Warriors’ 2015–2019 dominance made the NBA predictable and potentially drove fans away.
@@
Console vs PC 1. As a console gamer, I value its simplicity, exclusives, and plug-and-play convenience for relaxation.
2. As a PC enthusiast, I believe PCs provide unmatched customization, performance, and game variety for in-depth gaming.
@@
Working conditions in gaming 1. Crunch culture in game development is harmful and unsustainable; companies should prioritize employee wellbeing.
2. Crunch may not be ideal, but without it, could top games be made? Many vocations are stressful, not just gaming.
@@
Large vs indie game companies 1. Major game companies’ dominance stifles creativity; we need more space for indie developers to innovate.
2. Big companies like Nintendo prove that dominance doesn’t necessarily kill creativity, as they continually innovate and fund ambitious projects.
@@
eSports as a real sport 1. eSports, despite lacking physical exertion, is a sport due to its demand for strategy, teamwork, and skill.
2. Esports should be classified as competitive gaming, not traditional sports due to the lack of physical exertion.
@@
Objectification of female characters 1. I believe game developers overly sexualize female characters, neglecting character depth.
2. Video games are fantasy and idealized characters reflect artistic vision, not objectification.
@@
Copyright issues and modding 1. I support modders as they express themselves by reshaping games, while respecting authorship.
2. I support developers because unauthorized mods can threaten their control over their costly game creations.
Examples of viewpoints for r/NBA and r/gaming The table lists examples of controversial viewpoints for two non-political communities – r/Nba and r/gaming.
A.3. Identifying Relevant Subreddits
@@
To find our list of political and social issues-oriented subreddits, we use the following process. Starting with 2,040 subreddits from r/ListOfSubreddits, we look at the 50 hottest submissions in each subreddit as of January 5, 2024. We select subreddits where the majority of submissions are in English – classified using langdetect (Shuyo, 2010) – leaving us with 1,975 subreddits. Next, using GPT-3.5-Turbo we classify whether post titles are political (see below for prompt). We follow Piccardi et al. (2025) and use the definition of political content proposed by Pew Research Center (Bestvater et al., 2022). We choose subreddits where more than $80\%$ of submissions are labelled as being political (N=48). From the remaining subreddits, we select our final list of 23 based on subreddit size. In our final survey, participants also had the option to list an additional subreddit they were an active member of, bringing us to 33 subreddits in total.
@@
Political content on Reddit is varied and can be about officials and activists, social issues, or news and current events. Looking at the following post title, would you categorize it as POLITICAL or NOT POLITICAL content?
Answer 1 if it is POLITICAL, 0 otherwise.
Post: \${POST}
Answer:
A.4. Knowledge Quiz
@@
To validate that participants were members of the subreddits they selected, they were given a knowledge screening quiz. In the quiz, the participants were shown three post titles, two of which belonged to the top 50 hottest submissions of their selected subreddit and one that came from the top 50 hottest submissions of a different political subreddit. We manually checked that the title from the other subreddit was not relevant to the selected subreddit. The titles used in the knowledge quiz are listed in Table 7.
@@
Table 7. Examples of post titles used in the knowledge quizzes. Participants must correctly select the post titles that belong to the subreddit (i.e., those in columns “Title 1” and “Title 2”).
Post titles used in the knowledge quizzes The table shows post titles used in the knowledge quizzes for four subreddits — r/worldnews, r/socialism, r/politics, and r/unitedkingdom. There are three post titles for each subreddit.
A.5. Survey Details
@@
A.6. Descriptive Statistics
Finally, we provide the descriptive statistics of the variables we use in our linear mixed-effect models in Table 8 and the disaggregated content removal rates by subreddit (Fig. 8).
Table 8. An overview of descriptive statistics for all introduced variables (min, max, mean, median).
@@ A.6. Descriptive Statistics
Dependent Variables Share Likelihood 1 7 3.87 4
Upvote Likelihood 1 7 5.47 6
Controls WTSC 1.5 5.5 3.68 3.63
Agreement Intensity 1 3 2.17 2
Male (0/1) 0 1 0.73 1
White (0/1) 0 1 0.67 1
Democrat (0/1) 0 1 0.37 0
Account Tenure 0 16 6.16 6
Posting 1 4 2.53 2
@@
Inclusion 1 11 6.39 7
Diversity 1 11 6.33 6
Content Removal Rate (%) 15.99 41.46 30.63 34.47
Descriptive statistics for all study variables. The table reports minimum, maximum, mean, and median values for dependent, control, and independent variables.
Figure 8. Disaggregated content removal rates for the twelve subreddits included in our analyses. Content removal rates range from $16.0\%$ for r/conspiracy to $41.5\%$ for r/changemyview.
Bar chart of content removal rates across subreddits. Bar chart of content removal rates across subreddits. The figure ranks subreddits by the percentage of content removed by moderators. changemyview has the highest removal rate at 41.5%, followed by feminism (39.0%) and subredditdrama (38.7%). Lower rates are observed in economics (18.4%), latestagecapitalism (18.3%), and conspiracy (16.0%). Overall, the figure highlights substantial variation in moderation intensity across communities, with some subreddits removing over twice as much content as others.
@@ Appendix B Robustness Checks
In this section, we conduct robustness checks on our analyses. First, we present comparisons of our dataset post-filtering with the unfiltered dataset. Second, we show results using an alternative dichotomization of Incongruency. Then, we report using results with ModRatio for moderation. In our pre-registration, we planned on using both ModRatio and Content Removal Rate as measures of community moderation, but ultimately removed ModRatio as the two were correlated. Then, we provide results for our analysis, including a quadratic term for Content Removal Rate. Finally, we report results using an ordinal mixed-effects model.
B.1. Comparing filtered dataset
Variable Test Filtered Mean Unfiltered Mean Test Statistic p-value
Comment Likelihood Welch’s t-test 3.88 3.79 t(926.4)=-0.67 0.51
Upvote Likelihood Welch’s t-test 5.47 5.40 t(936.0)=-0.65 0.52
WTSC Welch’s t-test 3.68 3.68 t(934.7)=-0.01 0.99
Agreement Intensity Welch’s t-test 6.17 6.14 t(950.9)=-0.62 0.54
Gender Chi-squared $\chi^{2}(2)=0.01$ 0.99
White (0/1) Chi-squared $\chi^{2}(1)=0.00$ 0.98
Political Affiliation Chi-squared $\chi^{2}(4)=0.01$ 0.99
Account Tenure Welch’s t-test 6.16 5.78 t(928.2)=-1.63 0.10
Posting Welch’s t-test 2.53 2.51 t(948.5)=-0.40 0.69
Incongruency Chi-squared $\chi^{2}(1)=0.00$ 0.98
Diversity Welch’s t-test 5.33 5.31 t(957.6)=-0.19 0.85
Inclusion Welch’s t-test 5.39 5.37 t(948.6)=-0.14 0.89
Content Removal Rate Welch’s t-test 0.31 0.30 t(993.9)=-1.65 0.10
Table 9. We compare the filtered and unfiltered data, conducting t-test and $\chi^{2}$ differences for our independent and dependent variables. There are no statistically significant differences between our filtered and unfiltered data.
We compare the filtered dataset that we use in our analyses with the full, unfiltered dataset to examine if there are any participants or subreddits that are systematically excluded. As shown in Table 9, we conducted statistical analyses comparing the distributions of key variables between the full dataset and the filtered sample. Across all tests, we fail to reject the null hypothesis that filtered responses come from the same distribution as unfiltered responses. While there are no statistically significant differences between the filtered and unfiltered dataset, we do note the two following differences. First, participants in the filtered dataset have longer account tenures ($6.16$ years in the filtered dataset compared to $5.78$ for the unfiltered case). Second, subreddits in the filtered dataset have slightly higher content remove rates with a mean of $30.6\%$ compared to $29.8\%$, although this difference is not statistically significant (t(993.9)=1.65,p=0.10t(993.9)=-1.65,p=0.10).
B.2. Using an alternative operationalization of incongruency
Table 10. Our results are robust to alternative dichotomizations of opinion incongruency. We present the results of a linear mixed-effect model that predicts a participant’s likelihood of sharing their opinion to a subreddit using $\mathit{Incongruency}_{\mathrm{Alt}}$, which binarizes opinion incongruency based off a continuous score, as an independent variable and results after adding interaction effects.
$\mathit{Incongruency}_{\mathrm{Alt}}$ $\mathit{Incongruency}_{\mathrm{Alt}}$ w/ Interactions
Coef. SE $p$ Coef. SE $p$
Fixed Effects
(Intercept) $4.63$ $0.59$ $0.00^{***}$ $4.64$ $0.59$ $0.00^{***}$
@@ B.2. Using an alternative operationalization of incongruency
WTSC 0.28-0.28 $0.25$ $0.27$ 0.29-0.29 $0.25$ $0.25$
Agreement Intensity $0.50$ $0.06$ $0.00^{***}$ $0.50$ $0.06$ $0.00^{***}$
Male (0/1) 0.72-0.72 $0.55$ $0.19$ 0.72-0.72 $0.55$ $0.20$
White (0/1) $0.07$ $0.53$ $0.89$ $0.07$ $0.53$ $0.90$
Democrat (0/1) 0.71-0.71 $0.49$ $0.16$ 0.71-0.71 $0.50$ $0.16$
Account Tenure $0.08$ $0.24$ $0.74$ $0.09$ $0.24$ $0.71$
Posting $0.68$ $0.24$ $0.01^{**}$ $0.68$ $0.24$ $0.01^{**}$
@@
$\text{Incongruency}_{\mathrm{Alt}}$ (0/1) 1.21-1.21 $0.40$ $0.00^{**}$ 1.45-1.45 $0.70$ $0.04^{*}$
Inclusion $0.09$ $0.10$ $0.37$ $0.10$ $0.10$ $0.32$
Diversity 0.01-0.01 $0.10$ $0.93$ 0.04-0.04 $0.09$ $0.64$
Content Removal Rate 0.08-0.08 $0.08$ $0.35$ 0.06-0.06 $0.08$ $0.48$
$\text{Incongruency}_{\mathrm{Alt}}$ $\times$ Inclusion 0.53-0.53 $0.56$ $0.35$
$\text{Incongruency}_{\mathrm{Alt}}$ $\times$ Diversity $0.78$ $0.39$ $0.04^{*}$
$\text{Incongruency}_{\mathrm{Alt}}$ $\times$ Content Removal Rate 0.97-0.97 $0.51$ $0.06$
Model Fit
Marginal $R^{2}$ $0.232$ $0.236$
Conditional $R^{2}$ $0.798$ $0.797$
Note: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001
We demonstrate that our findings are robust to different definitions of our measurement for opinion incongruency. In Sec. 4.4, we define Incongruency as 1 if participants agree with the viewpoint and the majority of subreddit members are neutral or disagree; Incongruency is 0 if the participant agrees and the majority of the subreddit also agrees. In, our alternative dichotomization, $\mathit{Incongruency}_{\mathrm{Alt}}$, we compute the absolute difference between how much the participant agrees with a viewpoint, Agreement, and how much they believe the majority of subreddit members agree with the viewpoint. The continuous score ranges from 0 to $7$. Absolute differences greater than $4$ are coded as 1 (i.e., incongruent) and less than or equal to $4$ are coded as 0. The correlation (Pearson’s $r$) between Incongruency and $\mathit{Incongruency}_{\mathrm{Alt}}$ is $0.288$. Using Incongruency results in a slightly better model fit than $\mathit{Incongruency}_{\mathrm{Alt}}$, as indicated by its lower AIC score ($1462.7$ vs. $1489.2$).
As shown in Table 10, $\mathit{Incongruency}_{\mathrm{Alt}}$ has a negative relationship with the likelihood of sharing ($\beta=-1.21$, $p=0.002$). Furthermore, we observe qualitatively similar results with the interaction effects between $\mathit{Incongruency}_{\mathrm{Alt}}$ and Diversity ($\beta=0.78$, $p=0.044$) as well as $\mathit{Incongruency}_{\mathrm{Alt}}$ and Content Removal Rate ($\beta=-0.97$, $p=0.059$).
B.3. Replacing measures for moderation
Table 11. Including ModRatio does not substantially alter relationships between our independent variables and the likelihood of sharing a viewpoint. The results of a linear mixed-effect model that predicts a participant’s likelihood of sharing their opinion to a subreddit. We present results including ModRatio as an independent variable and results after adding interaction effects.